OpenStudy (anonymous):

Lyle has a fourteen-cup box of cereal. If one serving is three-fourths cup, how many servings are in the box?

OpenStudy (anonymous):

Well we know the 3/4 is less than 1 thus or answer should be more than 14 cups. If you a serving was 2 cups, how many servings are in the box then?

OpenStudy (anonymous):

ten and one-half eighteen and two-thirds fourteen and three-fourths nineteen and one-third

OpenStudy (anonymous):

\[\frac{ 14 }{ \frac{ 3 }{ 4 } }\] = \[\frac{ 14*4 }{ 3 } = 18+\frac{ 2 }{ 3 }\]